An agent consumes quantity (x1,x2) of goods 1 and 2. Here is his utility function: 𝑈(đ‘„1, đ‘„2) = âˆšđ‘„1 + 2 ∗ đ‘„2, his budget constraint is p1x1+p2x2 = m.

 a. Calculate the agent’s Marshallian demand (x*1 , x*2 ).

b. When would the agent’s consumer’s problem have a corner solution?


1
Expert's answer
2021-10-12T09:53:40-0400

Utility function is "x_1^3x_2" with budget constraint "p_1x_1+p_2x_2=m"

At optimum,  "\frac{MU

X_1}{

MU

X_2}



=\frac{p_1}{

p_2}"

a) Therefore,

"MUx_1=3x_1^2x_2\\MUx_2=x_1^3\\At\space optimum,\space \frac{3x_1^2x_2}{x_1^3}=\frac{p_1}{p_2}⇒\frac{3x^2}{x_1}=\frac{p_1}{p_2}\\⇒x_2=\frac{p_1}{3p_2}x1"

Putting value of x2 in x1, we get

"⇒p_1x_1+p_2(\frac{p_1}{3p_2})x1=m\\⇒p_1x_1+\frac{p_1}{3}x1=m\\⇒3p_1x_1+p_1x_1=3m\\⇒4p_1x_1=3m\\⇒x1*=\frac{3m}{4p_1}"

Similarly,

"x2^*=\frac{p_1}{3p_2}(\frac{3m}{4p_1})\\x2^*=\frac{m}{4p_2}"

Therefore, the agent's Marshallian demand "(x1^*,x2^*)= (\frac{3m}{4p_1},\frac{m}{4p_2})"

b) Hence, indirect utility function is derived by putting x1* and x2* in utility function is

"V(m,p1,p2)=(\frac{3m}{4p_1})^3(\frac{m}{4p_2})\\V(m,p1,p2)=\frac{3m^4}{16p_1p_2}"